OpenStudy (anonymous):

12ax^3+20bx^2+32cx what does this equal when factored out??

OpenStudy (anonymous):

What factors do all 3 terms have in common?

OpenStudy (anonymous):

i dotn know none of tis makes any sense because my teacher doesnt teach any of the material he just throws the assighnmkent and tests in your face and expect you to know what to do

OpenStudy (anonymous):

Ok, do you know what I mean by terms?

OpenStudy (anonymous):

How many terms does the expression you gave have?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Right. Ok so lets look at the first term \[12ax^3\] That's saying \(12 \times a \times x^3\) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok, so 12, a, and \(x^3\) are factors of \(12ax^3\). But some of those can be factored further. 12 for example can be factored into what?

OpenStudy (anonymous):

i dotn even know what factors are so yeah

OpenStudy (anonymous):

Ok. Factors are like terms, but terms are for addition and subtraction. Factors are for multiplication and division. When I say \(3 \times 4\) = 12, I'm saying that 12 is the PRODUCT of the FACTORS 4 and 3.

OpenStudy (anonymous):

So what are the different things you can multiply to get 12?

OpenStudy (anonymous):

4 and 3 and 1 and 12

OpenStudy (anonymous):

4*3, 1*12, and 6*2 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

We will ignore the factorization using 1's because you can have infinitely many 1's when you factor something, it's not very useful. 12*1 = 12, but 12*1*1*1*1*1 = 12 also.

OpenStudy (anonymous):

So lets look at the other 2 sets of factors. 4 and 3.

OpenStudy (anonymous):

Err 4*3 and 6*2

OpenStudy (anonymous):

Starting with 4*3. Now can either of those factors be factored further?

OpenStudy (anonymous):

Remember to ignore factorizations that use a 1.

OpenStudy (anonymous):

the 4

OpenStudy (anonymous):

And its factors are?

OpenStudy (anonymous):

2*2

OpenStudy (anonymous):

Right. So putting it all together we have that 12 = 2*2*3

OpenStudy (anonymous):

And if we factor the other way 12 = 6*2 = (3*2)*2 We get the same thing 2*2*3

OpenStudy (anonymous):

So no matter how you factor something you will end up with the same factors.

OpenStudy (anonymous):

So we can say that 12 has two factors of 2. And one factor of 3.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now then, what are the factors of \(x^2\)

OpenStudy (anonymous):

Remember that \(x^2\) = x*x

OpenStudy (anonymous):

x isnt a number so it doesnt have factors

OpenStudy (anonymous):

It is a number. We just don't know what number it is. So we can't factor further than x, but we do know that \(x^2\) has two factors of x.

OpenStudy (anonymous):

right? Because \(x*x = x^2\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok and a cannot be factored further cause we don't know what number it is either.

OpenStudy (anonymous):

Ok, so now putting it all together we have that \(12ax^2\) has two factors of 2, one factor of 3, one factor of a, and two factors of x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Whoops, it should be \(12ax^3\). So that's 3 factors of x

OpenStudy (anonymous):

So now you try to do the second term.

OpenStudy (anonymous):

2*10 2*5

OpenStudy (anonymous):

2*10 2*5

OpenStudy (anonymous):

20 = 2*10 = 2*(2*5) Is that what you mean?

OpenStudy (anonymous):

it is now

OpenStudy (anonymous):

Ok, so what are the factors of \(20bx^2\)

OpenStudy (anonymous):

list them just like I did for \(12ax^3\)

OpenStudy (anonymous):

2*5 b x*x

OpenStudy (anonymous):

You're missing an extra factor of 2.

OpenStudy (anonymous):

5*2 = 10, not 20

OpenStudy (anonymous):

5*2*2 = 20

OpenStudy (anonymous):

Try it again with the 3rd term just to be sure you have it down.

OpenStudy (anonymous):

i dont know anything for 32 other than 32 and 1

OpenStudy (anonymous):

Well it's an even number right?

OpenStudy (anonymous):

So you know at least that it has a factor of 2.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So what times 2 is 32?

OpenStudy (anonymous):

And if you aren't sure, how can you find out?

OpenStudy (anonymous):

If 2*h = 32 then h = 32/2 right?

OpenStudy (anonymous):

sorry my computetr is being extremely tarded and its pissing me off but its 16

OpenStudy (anonymous):

sorry my computetr is being extremely tarded and its pissing me off but its 16

OpenStudy (anonymous):

Ok right. And what are the factors of 16?

OpenStudy (anonymous):

2*8

OpenStudy (anonymous):

And the factors of 8?

OpenStudy (anonymous):

2*4

OpenStudy (anonymous):

And 4 can be factored

OpenStudy (anonymous):

2*2

OpenStudy (anonymous):

Ok let me show you a trick for keeping track of all this stuff.

OpenStudy (anonymous):

my computer keeps fluttering up and its pissing me off so i g2g

OpenStudy (anonymous):

ok